OpenStudy (anonymous):

The sum of x and 6 is equal to x multiplied by 3. Find the value of x.

OpenStudy (anonymous):

When we look at this problem, we see that we can deduce an equation off of what's given. When the problem says that the sum of x and 6, we can assume x+6. Also, when we see is equal to x multiplied by 3, we can also assume that means 3x. This knowledge gives us x+6=3x We can simplify this to be 6=2x This gives x=2

OpenStudy (anonymous):

so to solve for x divide each side by -2? so you would get, -2x = -6 \[\frac{ -2x }{ -2 } = \frac{ -6 }{ -2 }\] x=3?

OpenStudy (amigatour):

\[x+6=3x \] -3x-6 = -3x-6 -2x=-6=3 x=3

OpenStudy (anonymous):

yep 6/2 is 3. That's my bad